### v3

@@ 2.3 Comments
We collected 18.4 million comments on the videos (excluding comments on comments), associated with 13.5 million commenters. We collected these comments after the collection process was complete, in September 2024, 5 months after the day the videos were posted, using Ensemble Data  [^8] https://ensembledata.com/. The delayed comment collection process meant that we were not able to collect comments from posts that had been deleted since collection.
The median number of comments on the videos is 0, and the mean number is 4.92, with 51.9% of videos receiving zero comments.
When looking at the distribution of comment creation time depending on second of the minute and minute of the hour, we see a flat distribution (std dev. across each second of a minute: 652, std dev. across each minute of an hour: 3.7k, with a mean of 306k), indicating minimal scheduled posting of comments, and if automated comment posting is occurring, timing of posting is more sophisticated.
We also looked at average offset between video post and comment time, as shown in Figure 9. We see that for all commenters, the modal time to comment is just under 1 day, and that comments from more frequent commenters ($\geq 5$ comments in our dataset) have a modal response delay of 1 hour from post create time. We can also see that around 2/3 months after a video is posted, the commenting frequency on the video rapidly declines. This data shows us that while distribution processes on TikTok are fast, users are not engaging with the content until at least a few minutes past post time.
Looking at comment language, 47% of comments are unknown (46% of comments consist only of emojis, explaining this high percentage), 15% are in English, 9% are in Spanish, and 5% are in Arabic. The high percentage of emoji-only commenting on TikTok could imply large quantities of inauthentic activity, as emojis would be an easy way to boost a video’s engagement while requiring limited effort on behalf of the bot.
Figure 9: Time offset between video and comment creation by commenter frequency. We show offset distributions for authors who have at last 1 comment, authors who have at least 5 comments, and authors who have at least 25 comments, to show how commenting frequency affects commenting speed.
@@ 2.4 Error Analysis
We received several different types of error message for a requested video, and we report the most common error types, excluding the error message ‘item doesn’t exist’, in Table 2. We can see that an extremely large number of video posts are deleted, or privatized (status_self_see). We also gain insight into what seems to be the TikTok moderation process, with $\sim$7of videos unavailable at the time of collection due to a review stage (status_reviewing), $\sim$1of videos hidden after audit (status_audit_not_pass), and a small number potentially currently undergoing content classification for moderation (content_classification). We also see two error messages potentially related to geo-location, with cross_border_violation likely related to location-based content restrictions, for example when TikTok made new posts unavailable in Russia at the beginning of the 2022 invasion of Ukraine [27].
@@
Table 2: Distribution of most common post error messages, excluding ‘item doesn’t exist’. Reported percentage is out of all videos that at some point existed in our combined datasets. So from this we can infer that around 23% of videos on TikTok are deleted. Posts can have multiple error status messages, but there is no overlap between posts with status_self_see and status_deleted, so their similar count numbers seems to be a coincidence.
@@ 2.5 Inauthentic Activity
Given these constraints, we used simple heuristics for bot detection [32]. We find that there are 46 accounts with more than 100 videos each posted in the hour we looked at, and the top 5 highest frequency posting accounts, with between 1500 and 6000 videos in the hour we looked at, all have the default username of user followed by 9 random digits. Their videos have an average of 1.9 views each, but account for 0.28% of the total 1 hour dataset. Looking at comments, there are 831 accounts that have at least 10 comments and whose mean commenting frequency is higher than 1 every 3 seconds, with most of their comments (88%) consisting only of emojis. It’s clear from these initial statistics that there is automated activity on the platform, but these methods can only produce a rough lower bound on the extent of automated activity on the platform. We therefore hope future work can narrow our bounds on the amount of inauthentic activity on the platform and in this dataset.
@@ 3 Discussion
The data we report in this work is critical for calibrating our understanding of work on TikTok. Here we briefly illustrate the utility of the data with five exploratory case studies.
The invasion of Ukraine was touted as the ‘First TikTok War’ [33], and the way the invasion of Ukraine was covered on TikTok was investigated [34] as eyes turned towards the conflict. With this dataset, we can learn that while the average view count of posts from Ukraine (4.1k) was higher than the global average (2.2k), it is still lower than the average view count of posts from the US (6.3k), quantifying our understanding of its global attention on the day posts for this work were collected. We can also show that, in keeping with TikTok policy, no posts are tagged with a Russian origin (although of course workarounds such as VPNs for posting from Russia are possible).
In other places it’s been established that there is a tendancy in the global south towards engagement with content and culture from the global north [35]. Our data renders this in striking relief when we consider that the global south [36] accounts for 81% of video posts but only receive average view counts of 1.5k (compared to a baseline of 2.2k).
Finally, for those looking at topical issues on TikTok, our data provides an entrypoint by way of topical clusters. For work on Palestinian activism on TikTok [37], we can aid in initial investigations by showing that 410 videos in our hour dataset were clustered as a Palestine protest associated topic, but those videos had a significantly lower (p < 0.02) average view count of 440 views compared to the hour dataset view average of 2.5k. For work looking at the role of war on TikTok[38], we can show that 2,995 videos in our hour dataset were clustered as a military/soldier associated topic, and those videos received an average of 5.2k views. And for work looking at the link between TikTok and public health [39], we can add context by showing that 300 videos associated with the topic ‘Medical Examination and Healthcare Setting’ had a significantly higher (p < 0.02) average view mean of 28.8k views than the average.
@@
This dataset is not perfect. Collecting the data for this dataset took 5 months, meaning that the first batch of data was collected 2 months after the day chosen, and the last batch of data was collected 7 months after the day chosen. This introduces temporal bias into our dataset, where older videos will accumulate more engagement, and are also more likely to be taken down, or deleted. However, looking at our comment engagement offset (Fig. 6), we can see that the median time of engagement for all comment authors is around 29.9 hours after the video was posted. It seems likely that the median time from video post to comment engagement is similar to the median time offset for other engagement. This means that the vast majority of engagement with posts happens within two months, meaning that at least for engagement, the temporal bias of a collection process should have minimal effect. We also know that from looking at deletion rates at different minutes in our collection, there was an 8% percentage point increase in deletion rates from the first minute of our hour to the last minute.
@@ 4.1 ID Generation
Domingues et al. [45] found that the $3^{rd}$ lowest hexadecimal value determines the type of the ID (post, user, comment etc.), meaning the $52^{nd}$ to the $56^{th}$ bits [45].
We scraped a TikTok video metadata dataset of 225 thousand videos, attempting to capture videos that were likely produced in a variety of places to get high geographical distribution. The first method we used for this was datasets associated with accounts known to be in a specific country, which were available for Canada [46] and Germany [47]. For countries where we were unable to find country specific seedlists, we collected videos using the hashtag for the name of the country in the country’s official language. As long as a reasonable percentage of videos with the hashtag do indeed originate in that country, by looking at ID distribution, we can estimate whether a good ID coverage for that geographical region.
Country Method No. Videos Estimated Coverage
Canada Seedlist 118,994 99.93%
Germany Seedlist 35,256 99.94%
Nigeria Country Name Search 41,127 99.86%
Indonesia Country Name Search 19,509 99.83%
Brazil Country Name Search 14,079 99.77%
All 225,127 99.97%
Table 3: We estimate the percentage of videos we can capture, given the unique ID patterns we have found, using the Good-Turing estimator [48].
We then take the numerical IDs for those videos, and convert them to 64 bits of binary. We confirm that the first 32 bits of the ID correspond to the creation time of the video. This further implies that the TikTok ID generation system uses a Snowflake like method. We know that it isn’t exactly using the snowflake method, as it uses the first bit as part of the timestamp (Snowflake reserves this first bit as always zero) [41].
@@
The original Snowflake ID system contains a machine specific ID in the generated ID, where there are a finite number of machines that generate new IDs. In a global distributed system, it’s typically desired that different geographic regions (i.e. continents) are serviced by regional datacentres in order to reduce latency [49], and therefore machine IDs will likely be region specific. For that reason, we collected datasets correlated with globally distributed countries. We used Canada and Germany to cover North America and Europe as we had pre-existing, large datasets of videos from those countries. We searched for large country specific pre-existing datasets for other countries, but to the best of our knowledge finding an absence of such datasets, collected datasets by getting videos using the hashtag of the name of those countries, with the intuition that most videos featuring those hashtags will have been created in those countries. Due to the hashtag search mechanism being limited to 1000 videos, we then collected videos from the accounts discovered via the hashtag method to obtain more videos.
To ensure we have a representative dataset, we wanted to create some statistical assurances that we have captured enough of the possible ID patterns that we can say with confidence we are capturing the vast majority of videos. To do this, we can turn to the Good-Turing estimator [48], which, given our existing dataset of IDs, where we treat the non time encoding bits as the type of the ID, we can then estimate the number of each type of ID we have to estimate the probability of finding a previously undiscovered type of ID. This probability is then the probability of a video existing that we are not able to capture given our current technique. We report the corresponding percentage of coverage in Table 3.
@@
We then needed to generate these IDs, and request them from TikTok to see if they correspond to valid videos. To do this, we simply find all the possible variations of bit sequences for bits 42 to 63 (all the bits past the millisecond segment), and request each sequence of these for every millisecond. This results in 504 IDs to check for each millisecond.
@@
The Good-Turing estimator gives us confidence in what percentage of videos we have for each geographical area, but the question remains if we have good geographic coverage. After looking at the geographic distribution of the videos collected in the one minute from each hour dataset, we saw that the notable places where video counts were lower than others were East Asia, Oceania, India, and Russia. Though there are good reasons to believe that video counts in India and Russia would naturally be lower due to the ban of TikTok in those countries [24, 50], we collected extra datasets correlated with those regions (India, Australia, Japan, Russia) to check if there were previously undiscovered IDs, see details in Table 4. We found that the unique IDs we had previously collected covered at least 99% of the videos from those areas, and for India and Japan, at least 99.25%, which we feel is sufficient to ensure a dataset is representative of those areas.
Country Hashtag Name No. Videos Estimated Coverage Fetch Overlap
@@
Table 4: We estimate the percentage of videos we can capture, given the unique ID patterns we have found, using the Good-Turing estimator [48].
Specifically we notice that there are no video posts attributed to Russia in our dataset, which is notable given that both India and Russia have banned TikTok, but we still obtain posts attributed to India. To ensure that we have not completed missed ID aspects that are associated with Russia, we collected the full metadata of the 9% of posts we collected under the Russian hashtag that would not have been covered by our original fetch. We saw that none of them were attributed to Russia, only a range of other countries, implying that Russians using TikTok, or those ostensibly posting from Russia, are using VPNs, a common practice in other areas with banned internet services [51].
@@
We noticed that when we look at post time distribution over each second of the minute using the createTime attribute in the metadata, there is a 14% increase at the $0^{th}$ second compared to the rest of the seconds, but that this distribution is flat when looking at the create time of the post as derived from the ID. As detailed in Section 2.1.1, these videos at the $0_{th}$ second are significantly different from posts in other seconds. We hypothesize that this means that the create time in the ID is the time when a scheduled post is created, and the createTime attribute is the time when the video is made available as scheduled. This would imply that some of the posts we request are posts that are not yet available, and as such, would have an error message detailing that, which we hypothesize to be the status_self_see error message.
@@
We investigated the apparent counters found in the IDs, and trialled methods to predict the next valid video ID given previous counters. However, in the end we decided we wanted to get as complete a sample as we could, and therefore just requested all possible ID combinations. But if a method could be successfully developed here, it would greatly cut down the number of video IDs that would need to be requested, expanding the number of videos that could be collected through this method.
If we accepted a more probabilistic approach to video collection, we could have collected many more views for many fewer requests. However, this would have meant a less complete slice of TikTok. If from the data we have gathered, we can determine that ID patterns with more frequent video hits are not meaningfully different than ID patterns with less frequent occurrence of video posts, we could confidently roll out such a system. Figure 12 shows the percentage of videos we can expect to get by aiming for higher hit rates.
Figure 12: Success rate of ID generation vs percentage of videos successfully fetched out of all video we find for our combined dataset.
@@ 4.2 Collection
We collected two datasets, one of all video posts from 5-6pm UTC on the 10th April, 2024 (an arbitrary, mundane date, to obtain baseline data from the platform), and all videos from the $42^{nd}$ minute of each hour from 10pm UTC on the 9th April, till 10pm on the 10th April, to obtain across day representative data. The entire collection process took 5 months.
@@ 4.4.1 Topics
4.4.2 Classification of videos featuring children
We needed to build a classifier to classify each video in our 24 hour dataset as containing a child or not. The xclip video embeddings were used as the representations of each video. To train and test a classifier, we used a dataset of TikTok accounts [58]. The dataset contains 5,248 unique videos depicting children, including babies and young teenagers below the age of 13, and an additional 6,972 videos classified as not containing children. Videos were collected through a snowball sampling approach based on child-related keywords and are not restricted to a certain language or region. We evaluated methods via F1 on a 5-fold cross-validation train/test split. We tested several classifier methods, and found a stacking ensemble method using a logistic regression, histogram-based gradient boosting classifier  [^15] https://scikit-learn.org/stable/modules/generated/sklearn.ensemble.HistGradientBoostingClassifier.html, and an extra trees classifier [59] produced the best mean F1 of 0.86.
Given the difficulty of detecting children of varying age groups in contexts, the obtained F1 score of 0.86 seems particularly dependable. We looked at false positives and true negative predictions from the classifier, and found no general trend based on demographics. As stated by Schirmer et al. [58], the range of contexts in which videos feature children is vast: Videos contain vlogs of new mothers showing their newborn, parents playing with their toddlers, and families performing popular TikTok dances. Further, children might be shown only for a couple of seconds or throughout the whole video. Therefore, textual descriptions of videos can contain a lot of noise or might even miss child-related cues, making the performance of our classifier even more promising.

### v4

@@ 2.5 Comments
We collected 18.4 million comments on the posts (excluding comments on comments), associated with 13.5 million commenters. We collected these comments after the collection process was complete, in September 2024, 5 months after the day the posts were posted, using Ensemble Data  [^7] https://ensembledata.com/. The delayed comment collection process meant that we were not able to collect comments from posts that had been deleted since collection.
The median number of comments on the posts is 0, and the mean number is 4.92, with 51.9% of posts receiving zero comments.
When looking at the distribution of comment creation time depending on second of the minute and minute of the hour, we see a flat distribution (std dev. across each second of a minute: 652, std. dev. across each minute of an hour: 3.7k, with a mean of 306k), indicating minimal scheduled posting of comments, and if automated comment posting is occurring, timing of posting is more sophisticated.
We also looked at mean offset between post and comment time, as shown in Figure 10. We see that for all commenters, the modal time to comment is just under 1 day, and that comments from more frequent commenters ($\geq 5$ comments in our dataset) have a modal response delay of 1 hour from post create time. We can also see that around 2/3 months after a post is created, the commenting frequency on the post rapidly declines. This data shows us that while distribution processes on TikTok are fast, users are not engaging with the content until at least a few minutes past post time.
Looking at comment language, 47% of comments are unknown (46% of comments consist only of emojis, explaining this high percentage), 15% are in English, 9% are in Spanish, and 5% are in Arabic. The high percentage of emoji-only commenting on TikTok could imply large quantities of inauthentic activity, as emojis would be an easy way to boost a post’s engagement while requiring limited effort on behalf of the bot.
Figure 10: Time offset between post and comment creation by commenter frequency. We show offset distributions for authors who have at last 1 comment, authors who have at least 5 comments, and authors who have at least 25 comments, to show how commenting frequency affects commenting speed.
@@ 2.6 Error Analysis
We received several different types of error message for a requested video, and we report the most common error types, excluding the error message ‘item doesn’t exist’, here, and in Table 2. We can see that an extremely large number of video posts are deleted ($\sim$23%), or privatized (status_self_see) ($\sim$23%). We also gain insight into what seems to be the TikTok moderation process, with $\sim$7% of videos unavailable at the time of collection due to a review stage (status_reviewing), 1.18% of videos hidden after audit (status_audit_not_pass), and 0.16% currently undergoing content classification for moderation (content_classification). We also see two error messages potentially related to geo-location, with cross_border_violation (0.01%) likely related to location-based content restrictions, for example when TikTok made new posts unavailable in Russia at the beginning of the 2022 invasion of Ukraine [114].
@@
Table 2: Distribution of most common post error messages, excluding ‘item doesn’t exist’. Reported percentage is out of all posts that at some point existed in our combined datasets. So from this we can infer that around 23% of posts on TikTok are deleted. Posts can have multiple error status messages, but there is no overlap between posts with status_self_see and status_deleted, so their similar count numbers seems to be a coincidence.
@@ 2.7 Inauthentic Activity
Given these constraints, we used simple heuristics for bot detection [81]. We find that there are 46 accounts with more than 100 posts each posted in the hour we looked at, and the top 5 highest frequency posting accounts, with between 1500 and 6000 posts in the hour we looked at, all have the default username of user followed by 9 random digits. Their posts have an mean of 1.9 views each, but account for 0.28% of the total 1 hour dataset. Looking at comments, there are 831 accounts that have at least 10 comments and whose mean commenting frequency is higher than 1 every 3 seconds, with most of their comments (88%) consisting only of emojis. It’s clear from these initial statistics that there is automated activity on the platform, but these methods can only produce a rough lower bound on the extent of automated activity on the platform. We therefore hope future work can narrow our bounds on the amount of inauthentic activity on the platform and in this dataset.
@@ 3 Discussion
Altogether, the data in this work is critical for calibrating our understanding of TikTok, and ID sampling is critical to properly understand social platforms. We see substantial differences between outputs from our work, and the distributions available through the API [77] in Figures 1(c), 2(b), and 2(c), further validating the need for API audits [108]. We briefly illustrate the utility of the data with five exploratory case studies.
The invasion of Ukraine was touted as the ‘First TikTok War’ [74], and the way the invasion of Ukraine was covered on TikTok was investigated [111] as eyes turned towards the conflict. With this dataset, we can learn that while the mean view count of posts from Ukraine (4.1k) was higher than the global mean (2.2k), it is still lower than the mean view count of posts from the US (6.3k), quantifying our understanding of its global attention on the day posts for this work were collected. We can also show that, in keeping with TikTok policy, no posts are tagged with a Russian origin (although of course workarounds such as VPNs for posting from Russia are possible).
In other places it’s been established that there is a tendency in the global south towards engagement with content and culture from the global north [107]. Our data renders this in striking relief when we consider that the global south [112] accounts for 81% of posts but only receive mean view counts of 1.5k (compared to a baseline of 2.2k).
Finally, for those looking at topical issues on TikTok, our data provides an entry-point by way of topical clusters. For work on Palestinian activism on TikTok [72], we can aid in initial investigations by showing that 410 posts in our hour dataset were clustered as a Palestine protest associated topic, but those posts had a significantly lower (p < 0.02) mean view count of 440 views compared to the hour dataset view mean of 2.5k. For work looking at the role of war on TikTok[93], we can show that 2,995 posts in our hour dataset were clustered as a military/soldier associated topic, and those posts received an mean of 5.2k views. And for work looking at the link between TikTok and public health [95], we can add context by showing that 300 posts associated with the topic ‘Medical Examination and Healthcare Setting’ had a significantly higher (p < 0.02) mean view mean of 28.8k views than the mean.
@@
This dataset is not perfect. Collecting the data for this dataset took 5 months, meaning that the first batch of data was collected 2 months after the day chosen, and the last batch of data was collected 7 months after the day chosen. This introduces temporal bias into our dataset, where older posts will accumulate more engagement, and are also more likely to be taken down, or deleted. However, looking at our comment engagement offset (Fig. 6), we can see that the median time of engagement for all comment authors is around 29.9 hours after the video was posted. It seems likely that the median time from video post to comment engagement is similar to the median time offset for other engagement. This means that the vast majority of engagement with posts happens within two months, meaning that at least for engagement, the temporal bias of a collection process should have minimal effect. We also know that from looking at deletion rates at different minutes in our collection, there was an 8% percentage point increase in deletion rates from the first minute of our hour to the last minute.
@@ 4.1 ID Generation
[80] found that the $3^{rd}$ lowest hexadecimal value determines the type of the ID (post, user, comment etc.), meaning the $52^{nd}$ to the $56^{th}$ bits [80].
We scraped a TikTok video metadata dataset of 225 thousand posts, attempting to capture posts that were likely produced in a variety of places to get high geographical distribution. The first method we used for this was datasets associated with accounts known to be in a specific country, which were available for Canada [109] and Germany [69]. For countries where we were unable to find country specific seedlists, we collected posts using the hashtag for the name of the country in the country’s official language. As long as a reasonable percentage of posts with the hashtag do indeed originate in that country, by looking at ID distribution, we can estimate whether a good ID coverage for that geographical region.
We then take the numerical IDs for those posts, and convert them to 64 bits of binary. We confirm that the first 32 bits of the ID correspond to the creation time of the video. This further implies that the TikTok ID generation system uses a Snowflake like method. We know that it isn’t exactly using the snowflake method, as it uses the first bit as part of the timestamp (Snowflake reserves this first bit as always zero) [123].
@@
The original Snowflake ID system contains a machine specific ID in the generated ID, where there are a finite number of machines that generate new IDs. In a global distributed system, it’s typically desired that different geographic regions (i.e. continents) are serviced by regional datacentres in order to reduce latency [70], and therefore machine IDs will likely be region specific. For that reason, we collected datasets correlated with globally distributed countries. We used Canada and Germany to cover North America and Europe as we had pre-existing, large datasets of posts from those countries. We searched for large country specific pre-existing datasets for other countries, but to the best of our knowledge finding an absence of such datasets, collected datasets by getting posts using the hashtag of the name of those countries, with the intuition that most posts featuring those hashtags will have been created in those countries. Due to the hashtag search mechanism being limited to 1000 posts, we then collected posts from the accounts discovered via the hashtag method to obtain more posts.
To ensure we have a representative dataset, we wanted to create some statistical assurances that we have captured enough of the possible ID patterns that we can say with confidence we are capturing the vast majority of posts. To do this, we can turn to the Good-Turing estimator [84], which, given our existing dataset of IDs, where we treat the non time encoding bits as the type of the ID, we can then estimate the number of each type of ID we have to estimate the probability of finding a previously undiscovered type of ID. This probability is then the probability of a video existing that we are not able to capture given our current technique. We report the corresponding percentage of coverage in Table 3.
Country Method No. posts Estimated Coverage
Canada Seedlist 118,994 99.93%
Germany Seedlist 35,256 99.94%
Nigeria Country Name Search 41,127 99.86%
Indonesia Country Name Search 19,509 99.83%
Brazil Country Name Search 14,079 99.77%
All 225,127 99.97%
Table 3: We estimate the percentage of posts we can capture, given the unique ID patterns we have found, using the Good-Turing estimator [84].
@@
We then needed to generate these IDs, and request them from TikTok to see if they correspond to valid posts. To do this, we simply find all the possible variations of bit sequences for bits 42 to 63 (all the bits past the millisecond segment), and request each sequence of these for every millisecond. This results in 504 IDs to check for each millisecond.
@@
The Good-Turing estimator gives us confidence in what percentage of posts we have for each geographical area, but the question remains if we have good geographic coverage. After looking at the geographic distribution of the posts collected in the one minute from each hour dataset, we saw that the notable places where video counts were lower than others were East Asia, Oceania, India, and Russia. Though there are good reasons to believe that video counts in India and Russia would naturally be lower due to the ban of TikTok in those countries [118, 115], we collected extra datasets correlated with those regions (India, Australia, Japan, Russia) to check if there were previously undiscovered IDs, see details in Table 4. We found that the unique IDs we had previously collected covered at least 99% of the posts from those areas, and for India and Japan, at least 99.25%, which we feel is sufficient to ensure a dataset is representative of those areas.
Country Hashtag Name No. Posts Estimated Coverage Fetch Overlap
@@
Table 4: We estimate the percentage of posts we can capture, given the unique ID patterns we have found, using the Good-Turing estimator [84].
Specifically we notice that there are no posts attributed to Russia in our dataset, which is notable given that both India and Russia have banned TikTok, but we still obtain posts attributed to India. To ensure that we have not completed missed ID aspects that are associated with Russia, we collected the full metadata of the 9% of posts we collected under the Russian hashtag that would not have been covered by our original fetch. We saw that none of them were attributed to Russia, only a range of other countries, implying that Russians using TikTok, or those ostensibly posting from Russia, are using VPNs, a common practice in other areas with banned internet services [73].
@@
We noticed that when we look at post time distribution over each second of the minute using the createTime attribute in the metadata, there is a 14% increase at the $0^{th}$ second compared to the rest of the seconds, but that this distribution is flat when looking at the create time of the post as derived from the ID. As detailed in Section 2.1, these posts at the $0_{th}$ second are significantly different from posts in other seconds. We hypothesize that this means that the create time in the ID is the time when a scheduled post is created, and the createTime attribute is the time when the video is made available as scheduled. This would imply that some of the posts we request are posts that are not yet available, and as such, would have an error message detailing that, which we hypothesize to be the status_self_see error message.
@@
We investigated the apparent counters found in the IDs, and trialled methods to predict the next valid video ID given previous counters. However, in the end we decided we wanted to get as complete a sample as we could, and therefore just requested all possible ID combinations. But if a method could be successfully developed here, it would greatly cut down the number of video IDs that would need to be requested, expanding the number of posts that could be collected through this method.
If we accepted a more probabilistic approach to video collection, we could have collected many more views for many fewer requests. However, this would have meant a less complete slice of TikTok. If from the data we have gathered, we can determine that ID patterns with more frequent video hits are not meaningfully different than ID patterns with less frequent occurrence of posts, we could confidently roll out such a system. Figure 13 shows the percentage of posts we can expect to get by aiming for higher hit rates.
Figure 13: Success rate of ID generation vs percentage of posts successfully fetched out of all video we find for our combined dataset.
@@ 4.2 Collection
We collected two datasets, one of all posts from 5-6pm UTC on the 10th April, 2024 (an arbitrary, mundane date, to obtain baseline data from the platform), and all posts from the $42^{nd}$ minute of each hour from 10pm UTC on the 9th April, till 10pm on the 10th April, to obtain across day representative data. The entire collection process took 5 months.
@@ 4.4.1 Topics
4.4.2 Classification of posts featuring children
We needed to build a classifier to classify each video in our 24 hour dataset as containing a child or not. The xclip video embeddings were used as the representations of each video. To train and test a classifier, we used a dataset of TikTok accounts [117]. The dataset contains 5,248 unique posts depicting children, including babies and young teenagers below the age of 13, and an additional 6,972 posts classified as not containing children. Videos were collected through a snowball sampling approach based on child-related keywords and are not restricted to a certain language or region. We evaluated methods via F1 on a 5-fold cross-validation train/test split. We tested several classifier methods, and found a stacking ensemble method using a logistic regression, histogram-based gradient boosting classifier  [^14] https://scikit-learn.org/stable/modules/generated/sklearn.ensemble.HistGradientBoostingClassifier.html, and an extra trees classifier [85] produced the best mean F1 of 0.86.
Given the difficulty of detecting children of varying age groups in contexts, the obtained F1 score of 0.86 seems particularly dependable. We looked at false positives and true negative predictions from the classifier, and found no general trend based on demographics. As stated by [117], the range of contexts in which posts feature children is vast: Videos contain vlogs of new mothers showing their newborn, parents playing with their toddlers, and families performing popular TikTok dances. Further, children might be shown only for a couple of seconds or throughout the whole video. Therefore, textual descriptions of posts can contain a lot of noise or might even miss child-related cues, making the performance of our classifier even more promising.
When calculating the confidence intervals of our prediction, we wanted to account for the precision and recall of our classifier (0.88 and 0.81 respectively), the size of the test-set used to determine the accuracy of the classifier (6110), and how representative the test-set is to the real data. We quantify the representativeness of the test-set using a ‘data quality index’ $\rho$, a number between 0 and 1 [102]. We set this value to 0.5, given that the videos in the test-set are from a limited range of accounts. We then calculate $N_{eff}$, the effective sample size, via the design effect, $DEFF=\frac{1}{\rho^{2}},N_{eff}=\frac{N}{DEFF}$ [97]. We then bootstrap 2000 samples, first sampling our precision and recall given sampling variance $prec_{sample}\sim\beta(precN_{eff},(1-prec)N_{eff}),rec_{sample}\sim\beta(recN_{eff},(1-rec)N_{eff})$ [124]. We then account for systematic bias in the test-set by adding bias via a normal distribution $prec_{bias}\sim N(prec_{sample},0.2(1-\rho),rec_{bias}\sim N(rec_{bias},0.2(1-\rho))$ [122]. We use max $\pm 20\%$ here as that is the typical performance drop under distribution shift determined by [88]. Finally, we simulate the classification process in terms of how many examples are mis-classified.